OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

HOW do i do this one..?

OpenStudy (anonymous):

@some1 sorry to bother :c

OpenStudy (anonymous):

standard is ax+by=c

OpenStudy (anonymous):

1-2 ..5-(-2) <-- to find slope .right?

OpenStudy (anonymous):

yup. so -1/7=slope :)

OpenStudy (anonymous):

but fractions are hard... what i do next?

OpenStudy (anonymous):

do the same thing you've been doing. -1/7(-2)+2=b i don't have a calculator. O_O

OpenStudy (anonymous):

i do..

OpenStudy (anonymous):

dang i took forever..

OpenStudy (anonymous):

den you can do ittttt :) just convert -1/7 into decimal so it's easier to plug into your calculator. and ish okaiii

OpenStudy (anonymous):

16/7

OpenStudy (anonymous):

:D so.. x+y= 16/7

OpenStudy (anonymous):

YYYYYYYYYYYYYYYYYYYYYYEEEEEEEEEEEEEESHHHHHHHHHHHHHHHHHHHHHH =D

OpenStudy (anonymous):

bam!

OpenStudy (anonymous):

@some1 wrud?:)

OpenStudy (anonymous):

huge english project :/

OpenStudy (anonymous):

D: dang.. about what?

OpenStudy (anonymous):

omg thats not the right answer

OpenStudy (anonymous):

it's a project that basically portrays your personality in a book you make. You write about things that happen this year or about anything in your life, and you make a personal book about it. It's big...

OpenStudy (anonymous):

WAIT WHAT.

OpenStudy (anonymous):

its not right

OpenStudy (anonymous):

wait you wrote the slope wrong in the equation silly :P it's -1/7 * x+y=16/7

OpenStudy (anonymous):

sorry it's my fault i didn't see that. -.-

OpenStudy (anonymous):

wait.. in my answer thingy box

OpenStudy (anonymous):

have three spaces then y= then five spaces to fill in

OpenStudy (anonymous):

but you cant put these ---> /

OpenStudy (anonymous):

so try decimals.

OpenStudy (anonymous):

i have one space before i put x+y... it has to be 1 number..

OpenStudy (anonymous):

hm. check the coordinates again..did u post them right? cuz there seems to be no problem with the answer here...

OpenStudy (anonymous):

:/ the coordinates are right

OpenStudy (anonymous):

;asldkfj;aosidjf;lk what are we doing wrong?!?!?!?!

OpenStudy (anonymous):

idk :/oh well

OpenStudy (anonymous):

i'm sorry :'(

OpenStudy (anonymous):

its ok:)

OpenStudy (anonymous):

the answer was x+7y=6